a. Assume that a sample is used to estimate a population proportion p. Find the margin of error M.E. that corresponds to a sample of size 343 with 292 successes at a confidence level of 99.8%.

M.E.=

b. You measure 46 textbooks' weights and find they have a mean weight of 79 ounces. Assume the population standard deviation is 7.5 ounces. Based on this, construct a 90% confidence interval for the true population mean textbook weight.

Give your answers as decimals, to two places

< μ <

c.  Suppose n=40,¯x=38n=40,x¯=38 and σ=2.5σ=2.5. Based on this, what is the maximal margin of error associated with a 99% confidence interval for the true population mean.

Margin of error =

d. You measure 33 turtles' weights, and find they have a mean weight of 30 ounces. Assume the population standard deviation is 9.1 ounces. Based on this, construct a 99% confidence interval for the true population mean turtle weight.

Give your answers as decimals, to two places

± ounces
Answer #1

-Ans ☺ Given that 292. 2 292 n D. .851 343 1- 1-0,852 -0,148 - At 99,8-1. Confidence level the zin do 16.998 212 0.1002/2 = -2 - 3,090*0,01917 Es 0.059 Margin of error (M.E) = -0.059. b) M = 79 715 n 46 - ) 901. Confidence level is 1-0.90 ol 2/2 = GbAt 90%. Confidence intercal in 3 -Е 2-е 1-р 79-1.8191.2 M279+1,8191 77.18 L M L 80.81 © Girren that - point estimate Sample m= 2.576 0.3953 1.02 Norgen of erlor - E - 1102 cd, Meanla) - 30. - Standard deviatim (o)2911 Sample Size (n) 233 = 99-1. Conf